# Hard Negative Contrastive Learning for Fine-Grained Geometric Understanding in Large Multimodal Models

### v1

@@ 4.1.3 Overall Results
As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-MATH benchmarks. On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B. As Chimera-Reasoner-8B was trained on GeoQA Peng et al. (2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset. The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach. For the We-Math benchmark, MMGeoLM-Mammoth2-7B underperforms GPT-4o, Claude-3.5-Sonnet, and the open-source Qwen2.5-VL-7B-Instruct model. We attribute this to We-Math’s emphasis on recognition from visual elements rather than geometric mathematical reasoning, which limits the effectiveness of MMGeoLM-Mammoth2-7B’s geometric reasoning enhancement strategies. On the open-ended MM-MATH benchmark, MMGeoLM-Qwen2.5-7B performs poorly on hard problems, with less than 10% accuracy, but achieves over 55% accuracy on easy problems. Easy problems require fewer reasoning steps, allowing geometric element recognition to significantly enhance accuracy. In contrast, hard problems involve multi-step reasoning, where its impact is more limited. Compared to using the original AltCLIP and without MMCLIP (AltCLIP trained only with in-batch learning on MAVIS), the vision encoder trained with our constructed hard negative samples and the MMCLIP training method significantly enhances MMGeoLM’s geometric understanding capabilities.
Figure 3: Performance evaluation of MMGeoLMs built with six AltCLIP trained on different contrastive datasets.
Figure 4: Comparison of two image negative strategies: Mavis uses 100K samples from the Mavis dataset, while MM-MATH-Align uses our 4K constructed samples.
Figure 5: Retrieval performance of AltCLIP variants trained with different strategies. ‘Random + X’ means sequential training with random negatives followed by negatives created with the X strategy.
4.2 Experiments on Vision Encoder
We further quantitatively evaluate the performance of various vision encoders trained with different datasets and strategies. The experiments are divided into three parts: (1) performance evaluation of six LMMs trained with different AltCLIP variants across two benchmarks, (2) top-1 positive example retrieval performance evaluation of six AltCLIP models, and (3) varying number of hard caption negatives.
4.2.1 Setup
We compare the following six types of AltCLIP models are as follows:
  • Original AltCLIP: No additional data training.
  • Random AltCLIP: Trained with randomly selected negative samples within a batch (batch size = 10) using 400K MAVIS data.
  • Retrieved AltCLIP: Trained with 100K samples, each with 10 retrieval-based caption negatives.
  • Rule-based AltCLIP: Trained with 100K samples, each with 10 rule-based caption negatives.
  • Negative-images AltCLIP: Trained with 4K samples, each with 10 hard image negatives.
  • Retrieved + Rule-based + Negative-images: Trained on all three datasets combined.
The Random AltCLIP follows OpenAI’s training approach Radford et al. (2021), while the Retrieved, Rule-based, Negatives-images and Retrieved + Rule-based + Negatives-images AltCLIP models employ our proposed MMCLIP training approach. In the second and third stages, LMMs are trained on 67K G-LLaVA alignment data and 17K open-ended geometric reasoning problems.
For the retrieval experiments, we assess the retrieval performance of trained AltCLIP models by selecting positive samples from a large pool of negative samples. Retrieval performance is evaluated using the Hit@1 metric. We construct four types of negative datasets: Random Negative: Caption negatives randomly selected in a training batch. Retrieval Negative: Caption negatives retrieved using the SimANS Zhou et al. (2022) model, selecting the top 100 most similar captions to the positive sample. Rule-based Negative: Caption negatives generated according to the rules defined in section 2.2. Image Negative: Image negatives generated from real-world geometric problems by MM-MATH.
For the number of negative experiments, we investigate the impact of hard caption negative samples on MMGeoLM varying from 5 to 50.
4.2.2 Results
As shown in Figure 4, LMMs trained with hard negative samples consistently improve performance on the GeoQA and MM-MATH benchmarks. Notably, on MM-MATH, Negative-images AltCLIP, trained with only 4K image negative samples, surpasses Retrieved AltCLIP and Rule-based AltCLIP by 0.2% and 1.7%, respectively. On MM-MATH, Rule-based, Retrieved, and Image-negative AltCLIP consistently outperform Random AltCLIP on easy and medium-level questions. These results confirm the effectiveness of both hard negative construction methods. Furthermore, AltCLIP trained with a combination of caption and image negatives achieves the best performance on both GeoQA and MM-MATH, demonstrating their complementary effect in improving model performance. Further experiments on image negatives (Figure 4) reveal that Negative-image-MM-MATH-Align, constructed from real-world images, outperforms Negative-image-Mavis. This suggests that high-quality, realistic image negatives have a more pronounced impact on model performance than large-scale synthetic counterparts.
The results of the retrieval performance experiments are shown in Figure 5. Under the negative-images setting, Negative-images AltCLIP achieves a retrieval score of 23%, while other AltCLIP variants remain below 15%. These results highlight the difficulty of retrieving realistic images. Notably, the dataset includes only 4,021 image negative samples, which is likely insufficient to fully exploit the model’s performance potential. In the retrieval-based and rule-based negative settings, Random AltCLIP outperforms Rule-based AltCLIP in retrieval-based settings and exceeds Retrieved AltCLIP in rule-based settings. These results indicate that while hard negative training enhances performance within specific domains, it may limit generalization. Furthermore, pre-training on random negatives followed by hard negatives improves both generalization and domain-specific performance, ultimately enhancing overall model effectiveness.
@@ 4.2.2 Results
Figure 6: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative samples.
@@
As shown in Figure 6, MMGeoLM’s performance on MM-MATH improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point. These results indicate that hard caption negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance. Training with only five hard negative samples outperforms training with ten randomly selected negatives (as shown in Table 1), demonstrating that hard negatives enhance multimodal mathematical reasoning more effectively.
5 Related Work
@@
Mathematics-related research has recently received significant attention in LLMs and LMMs. For text-only mathematical reasoning, several works employ external tools like Tora Gou et al. (2023) or intermediate step decomposition methods such as MAmmoTH Yue et al. (2023), Metamath Yu et al. (2023) and Math-Shepherd Yue et al. (2023). For multimodal mathematics tasks involving images, benchmarks have been proposed, including mathematical competition-oriented OlympiadBench He et al. (2024), geometry-focused datasets like Geometry3K Lu et al. (2021), VisScience Jiang et al. (2024), UniGeo Chen et al. (2022a) and GeoQA Chen et al. (2022a). Many works take effort on training multimodel math, including G-llava Gao et al. (2023), Meta-LLaVA Shi et al. (2024) and MAVIS Zhang et al. (2024c). However, these multimodal approaches typically suffer from inaccuracies in recognizing geometric elements, where our paper specifically addresses the enhancement of visual recognition accuracy.
@@
Vision encoders play a crucial role in understanding image content. While previous works Zhang et al. (2024a); Doveh et al. (2023a, b); Singh et al. (2023) have leveraged negative captions to enhance image comprehension, but not address the recognition of geometric elements within images. Methods such as NegCLIP Yuksekgonul et al. (2022) and TriCLIP Yang et al. (2024) focus on learning from negative samples, but their applicability is limited by the constrained number of negative samples. In this work, we propose a robust negative learning strategy that explicitly constructs negative samples for geometric elements while enabling arbitrary expansion, thereby improving geometric understanding in multimodal tasks.
6 Conclusion
@@
This paper introduced MMCLIP, a novel hard negative contrastive learning framework to enhance LMMs’ geometric reasoning by refining their vision encoders with systematically generated image and text-based negatives. Our resulting model, MMGeoLM, significantly outperformed existing models, even surpassing GPT-4o on key benchmarks. This work demonstrates the power of tailored hard negative sampling for developing LMMs with fine-grained visual understanding in specialized domains such as geometric understanding and reasoning.
Limitations
Despite the effectiveness of the proposed image hard negatives constructed by perturbing Python scripts, the method heavily relies on the accuracy of LLM-generated code and captions. Any systematic biases in these synthetic constructions may introduce artifacts that differ from human-designed geometric problems. As a result, the model’s performance on real-world visual inputs remains uncertain and requires further validation on diverse real-world datasets.
Appendix A Prompts
After multiple experimental iterations, we constructed two different types of prompts for our experiments, as illustrated in Figures 7 and 8. Specifically, Figure 7 demonstrates the prompt constructed by inputting the textual description and answer of realistic geometry problems to obtain the corresponding geometric diagram Python script. Meanwhile, Figure 8 illustrates the prompt obtained by modifying captions from positive examples to create slightly perturbed negative samples.
Figure 7: Python script prompt from question and true answer in geometric problems
Figure 8: Negative caption generation prompt from positive captions
Appendix B Figure Comparison
Figure 9 presents illustrative examples of image negative sample, covering analytical geometry and planar geometry. Geometric elements include triangles, quadrilaterals, as well as various relational properties such as perpendicularity and intersection. Additionally, certain metric properties absent from the original diagrams are explicitly indicated. Specifically, the first column shows the original diagrams, and the second column depicts diagrams generated from the given questions and the true answer, which, despite slight discrepancies, accurately capture the overall outlines of the original diagrams. The third and fourth columns display the constructed negative diagram instances—generated based upon their corresponding positive examples—demonstrating notable similarities but also clear and meaningful distinctions.
Figure 9: Comparison of generated geometric figures. The first column shows the original images, the second column presents the positive samples, and the last two columns illustrate the constructed negative samples.
Appendix C Case Study
Figure 10: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. MMGeoLM correctly answers the problem using a different solution compared to True Answer, while GPT-4o generates a non-existent element $\triangle ABD$.
Figure 11: Comparison of GPT-4o and MMGeoLM in geometric problem-solving. Both models produce incorrect answers, but MMGeoLM’s solution is closer to the True Answer.
C.1 Case Study
In Figures 10 and 11, we compare the solutions generated by GPT-4o and our MMGeoLM. In Figure 10, the red text highlights GPT-4o’s misrecognition of two triangles in the image, leading to an incorrect final result. In contrast, MMGeoLM produces a correct solution, albeit through a reasoning process different from the ground-truth answer. This highlights MMGeoLM’s capability to generate diverse problem-solving approaches. Figure 11 illustrates errors in image-based reasoning for both MMGeoLM and GPT-4o. While both models make mistakes, MMGeoLM’s reasoning aligns more closely with the correct solution, demonstrating improved mathematical reasoning. Overall, our trained MMGeoLM model demonstrates greater reasoning diversity and enhanced geometric problem-solving performance compared to current mainstream models.

### v2

@@ Introduction
As shown in Table 2, our proposed MMGeoLM achieves state-of-the-art performance on the MathVista and MM-Math benchmarks. On the GeoQA benchmark, MMGeoLM-Qwen2.5-7B lags 0.4% behind Chimera-Reasoner-8B. As Chimera-Reasoner-8B was trained on GeoQA (Peng et al. 2024b), MMGeoLM-Qwen2.5-7B achieves the best performance among other models that were not trained on this dataset. The improved performance in geometric problem-solving demonstrates the effectiveness of our training approach. For the We-Math benchmark, MMGeoLM-MAmmoTH2-7B underperforms GPT-4o, Claude-3.5-Sonnet, and the open-source Qwen2.5-VL-7B-Instruct model. We attribute this to We-Math’s emphasis on recognition from visual elements rather than geometric mathematical reasoning, which limits the effectiveness of MMGeoLM-MAmmoTH2-7B’s geometric reasoning enhancement strategies. On the open-ended MM-Math benchmark, MMGeoLM-Qwen2.5-7B performs poorly on hard problems, with less than 10% accuracy, but achieves over 55% accuracy on easy problems. Easy problems require fewer reasoning steps, allowing geometric element recognition to significantly enhance accuracy. In contrast, hard problems involve multi-step reasoning, where its impact is more limited. Compared to using the w/o MMCLIP (AltCLIP trained only with random negatives in a batch), the vision encoder trained with our constructed hard negative samples significantly enhances MMGeoLM’s geometric understanding capabilities.
Ablation Study 1: Different Negatives Types
Vision Encoder #Num. GeoQA $M$-MATH
Easy Med Hard Avg
Original - 45.4 40.5 20.8 4.5 23.8
$\text{Random}_{\text{10}}$ 400K 45.0 39.8 22.7 4.5 24.9
$\text{Retrieval}_{\text{10}}$ 100K 53.0 43.6 24.8 4.5 26.6
$\text{Rule}_{\text{10}}$ 100K 48.8 45.4 27.6 4.5 28.1
$\text{Image-based}_{\text{10}}$ 4K 54.9 45.6 26.5 4.5 29.0
$\text{Retrieval}_{\text{10}}$$\text{Rule}_{\text{10}}$ 200K 56.5 47.5 28.3 4.5 29.4
All Negatives 204K 58.2 49.8 29.1 4.5 30.4
Table 3: Ablation study on vision encoder training with different types and quantities of hard negatives. The bottom panel reports performance when combining multiple types.
To further evaluate the performance of vision encoders trained with different types of hard negatives, we compare seven negatives summarized in Table 3. All models are based on AltCLIP. The Original vision encoder is directly initialized without any additional training, while $\text{Random}_{\text{10}}$ follows the conventional in-batch random negatives sampling strategy (Radford et al. 2021). The remaining five vision encoders are trained with our proposed MMCLIP method. The subscript 10 indicates a 1:10 positive-to-negative ratio.
Results
As shown in Table 3, vision encoders trained with hard negatives consistently outperform random in-batch negatives. even with only 4K image-based negatives , the $\text{Image-based}_{\text{10}}$ encoder achieves the highest average accuracy (29.0%) among all single hard negative methods on MM-Math. Furthermore, while $\text{Retrieval}_{10}$ and $\text{Rule}_{10}$ negatives individually provide moderate gains, their combination further improves performance (29.4%), indicating their complementary nature. The best results of All Negatives are achieved when all three types of negatives are combined, demonstrating that diverse hard negatives collectively enhance the model’s geometric understanding. These findings underscore the importance of both the type and diversity of hard negatives in contrastive training for vision encoders.
Ablation Study 2: Image-based Negatives
Method #Num. GeoQA $M$-Math
Easy Med Hard Avg
Rule-Img 100K 52.9 44.0 25.4 1.0 28.3
Perturb-Img 4K 54.9 45.6 26.5 4.5 29.0
$\Delta$ (Perturb-Rule) -96K +2.0 +1.6 +1.1 +3.5 +0.7
Table 4: Effect of different image negative construction strategies. $\Delta$ rows give absolute gains of Perturb-Img over Rule-Img.
2.1. Code Perturbation vs. Rule-design: Assessing Image Negatives Construction Methods
We contrast two image–based negative construction methods. Perturb-Img applies our code perturbation method to 4K images, generating negative images that preserve similar structure while introducing subtle differences from the original diagrams. In contrast, Rule-Img is built from manually designed templates in the MAVIS corpus, enabling large-scale expansion. We construct 100K samples and retrieve most similar images as hard negatives. To ensure a fair comparison, we fix the positive-to-negative ratio at 1:10 in both settings. Results in Table 4 show that the 4K Perturb-Img setting exceeds the 100K Rule-Img on both GeoQA (54.9 vs. 52.9) and MM-Math (29.0 vs. 28.3). These results highlight that code-perturbed negatives, despite being smaller in scale, outperform synthetic data due to their semantic diversity and closer resemblance to real-world tasks.
2.2. Length/Angle Marked Matters? Cross-Modal Consistency Analysis
Figure 4: An example of vision encoder training data with or without numeric markings.
Visual Marking GeoQA $M$-MATH
Easy Med Hard Avg
w/ length/angle/etc. 54.9 45.6 26.5 4.5 29.0
w/o length/angle/etc. 51.7 50.0 25.1 4.5 28.6
Table 5: Impact of numeric markings on vision encoder training. w/ and w/o denote settings with and without such numeric annotations, respectively.
[…]
2.3. Hard Negative Performance Gains or Trade-off for In-Batch Training
Negative Strategy $M$-MATH
Easy Med Hard Avg
In-Batch 41.8 23.7 4.5 25.5
+image-based NS 51.7 (+9.9↑) 29.0 (+5.3↑) 4.5 32.1 (+6.6↑)
Table 6: Comparison of In-Batch and In-Batch training with the proposed image-based negative strategy. Mean score over 5 random seeds on 4K image-based negatives.
We leverage the proposed image-based negative strategy for the in-batch training to verify the effectiveness. In the in-batch negatives setting, each sample in the batch serves as a hard negative for others, with traditional in-batch typically selected at random. For a fair comparison, both methods use 10 negative examples per positive sample. As shown in Table 6, in-batch training with our proposed image-based negative strategy outperforms In-Batch on the easy, medium, and average subsets under the same amount of training data. This highlights the advantage of explicitly constructed image-based negatives in guiding the model to better distinguish visually similar examples, thus achieving higher overall accuracy. Furthermore, these results indicate that carefully selected mutual hard negatives within batches lead to more effective contrastive learning.
Ablation Study 3: Training Strategy and Robustness
@@ 2.3. Hard Negative Performance Gains or Trade-off for In-Batch Training
Figure 5: MMGeoLM Performance with AltCLIPs trained with varying numbers of hard negative ratio.
3.1. Data Scaling: More Hard Negatives Are Not Always Better
@@
Our proposed MMGeoLM method allows flexible scaling of negatives. However, whether performance improves proportionally with an increasing number of hard negatives remains unclear. To address this, we investigate the impact of varying the number of retrieval-based negatives per sample from 5 to 50. As shown in Figure 5, MMGeoLM’s performance on MM-Math improves with increasing hard negative samples from 5 to 30, but slightly declines beyond this point. These results indicate that hard negative samples have diminishing returns beyond a certain threshold, with excessive examples reducing model performance.
3.2. Robustness to Geometric Image Transformations
Transformers Negatives $M$-MATH
Easy Med Hard Avg
No-Changing Text-based 48.6 29.1 9.0 31.9
[…]
Downscaling Text-based 45.6 26.8 0.0 28.4
No-Changing Image-based 45.6 26.5 4.5 29.0
[…]
Downscaling Image-based 48.2 25.6 4.5 28.7
Table 7: Robustness evaluation of MMGeoLM under image transformers (rotation and downscaling).
Diagrams in real exams frequently undergo transformations, such as rotation and scaling. To evaluate MMGeoLM’s robustness to these variations, we apply two image perturbations: (1) a $30^{\circ}$ clockwise rotation, and (2) downscaling images by a factor of 0.5. Table 7 summarizes the results under both text-based and image-based negative conditions. Model performance remains largely stable under a $30^{\circ}$ rotation, showing only a minor accuracy drop of 0.9 (31.9 $\rightarrow$ 31.0) in the text-based setting. This suggests the model has adequate robustness against moderate image rotations. In contrast, downscaling significantly impacts accuracy, dropping text-based performance from 31.9 to 28.4. This decline indicates that textual captions are particularly sensitive to reduced image resolution whereas the image-based negative demonstrates stronger robustness, benefiting from diagram scale variations.
Related Work
@@ 3.2. Robustness to Geometric Image Transformations
Mathematics-related research has recently received significant attention in LLMs and LMMs. For text-only mathematical reasoning, several works employ external tools like Tora (Gou et al. 2023) or intermediate step decomposition methods such as MAmmoTH (Yue et al. 2023), Metamath (Yu et al. 2023) and Math-Shepherd (Yue et al. 2023). Many multimodal mathematics benchmarks have been proposed, including mathematical competition-oriented OlympiadBench (He et al. 2024), geometry-focused datasets like Geometry3K (Lu et al. 2021), VisScience (Jiang et al. 2024), UniGeo (Chen et al. 2022a) and GeoQA (Chen et al. 2022a). Many works take effort on training multimodel math, including G-llava (Gao et al. 2023), Meta-LLaVA (Shi et al. 2024) and MAVIS (Zhang et al. 2024b). However, these models often struggle with fine-grained recognition of geometric elements.
@@
Vision encoders are critical for capturing spatial and structural information in multimodal tasks. Prior works (Zhang, Awal, and Agrawal 2024; Doveh et al. 2023a, b; Singh et al. 2023) enhance image understanding using negative captions, but rarely target fine-grained geometric recognition. Methods such as NegCLIP (Yuksekgonul et al. 2022) and TriCLIP (Yang et al. 2024) focus on learning from negative samples, yet their effectiveness is limited by small or randomly selected negatives. In this work, we propose a scalable and geometry-aware negative learning strategy that explicitly constructs task-relevant hard negatives to improve geometric understanding.
Conclusion
@@
This paper introduced two types of hard negatives, image-based and text-based negatives, targeted at geometric element understanding to enhance LMMs’ geometric reasoning. Our resulting model, MMGeoLM, significantly outperformed existing models, even surpassing GPT-4o on key benchmarks. This work demonstrates the power of hard negative contrastive learning for improving LMMs with fine-grained visual understanding. Extensive ablation studies confirm the effectiveness of hard negatives, highlighting the crucial role advancing in multimodal mathematical reasoning.